OpenStudy (anonymous):

a1=32, r=-0.06 Find the sum of the given infinite geometric series. I hate sums :(

OpenStudy (anonymous):

You're given the infinite geometric series \[\sum_{n=0}^\infty32\left(-0.06\right)^n\\ \sum_{n=0}^\infty32\left(-\frac{3}{50}\right)^n\] The sum of an infinite geometric series is \[\sum_{n=0}^{\infty}ar^n=\frac{a}{1-r}\]